The following ANOVA table is from an experiment that had a group receiving a placebo as the no treatment control, a group that received a dose of experimental compound A, another group receiving compound B while a final group received both drugs in combination. Each of the groups contained 10 subjects. The following information has been calculated but important pieces are missing.

Source of Variation Sum of Squares df Mean squares F-ratio P
Groups (treatments) 8.804 0.004
Error 10.512
Total 19.316

Given the above table calculate the degrees of freedom (df) column.

We need the df only. please show your work and I will rate your answer for sure. Thank you
Answer #1

Total no. of treatments=4 (1. a placebo 2. Treatment A, 3. Treatment B, 4. Interaction AB )

Total no. of observations=10*4=40